OpenStudy (anonymous):

3/7 + 2/9 = 5/16 is this right and in the lowest terms

OpenStudy (asnaseer):

you answer is not right. can you please show how you arrived at that answer so that I can help you spot where you may have gone wrong?

OpenStudy (anonymous):

\[\frac{3}{7}+\frac{2}{9}=\frac{41}{63} \]

OpenStudy (anonymous):

\[\frac{a}{b}+\frac{c}{d}=\frac{a d+b c}{b d} \]

OpenStudy (anonymous):

Easiest way to do this problem? Convert these fractions to have the same number on the bottom then add the two top numbers and put it over the number you converted it into then reduce.